# S4ECG: Exploring the impact of long-range interactions for arrhythmia prediction

### v1

@@ 4.2 Clinical implications
Our findings demonstrate substantial clinical implications that address critical gaps in contemporary arrhythmia monitoring and management. The demonstration that moderate temporal windows (10-20 minutes) yield optimal performance fundamentally challenges conventional arrhythmia detection paradigms.
@@
Furthermore, the superior out-of-distribution performance of multi-epoch models indicates a high degree of robustness—a critical requirement for real-world deployment, which was explicitly acknowledged for example in the FDA’s Software as a Medical Device (SaMD) guideline [26] and the action plan on AI/ML-based SaMD [27]. This robustness addresses the well-documented challenge of domain shift in medical AI applications, where models frequently underperform when applied to patient populations, recording devices, or clinical environments that differ from training conditions. The enhanced generalizability of our approach suggests improved performance across diverse healthcare settings, patient demographics, and ECG acquisition systems without requiring extensive retraining or calibration. More specifically, we anticipate impact on the following use cases:
@@
Enhanced diagnostic precision and clinical burden reduction Improved overall predictive performance at fixed sensitivity can translate into higher specificity, reducing false positives, unnecessary clinical interventions, costs, and patient anxiety. False positive rates have been reported to be high in certain conventional automated systems [22, 28], contributing to increased emergency department utilization from consumer-grade devices [29, 30]. As a result, the proposed approach can improve patient experience and resource utilization without compromising diagnostic safety.
@@
Advanced arrhythmia characterization and temporal dynamics Understanding the temporal dynamics characteristics of arrhythmia is key for more fine-grained understanding of arrhythmia subclasses, such as paroxysmal and persistent in the case of atrial fibrillation, and is believed to lead to clinically actionable insights into disease progression and treatment response [31, 32]. The improved accuracy but also temporal consistency, see Figure 3, of the proposed approach aligns with this goal.
@@
Atrial fibrillation burden assessment Precise AF burden quantification is increasingly recognized as a critical determinant of stroke risk and therapeutic decision-making [33]. Studies demonstrate that even modest AF burdens (>>0.5%) correlate with increased thromboembolic risk, emphasizing the clinical importance of accurate measurement. Our approach enables continuous, high-resolution burden assessment that can inform both acute and chronic management strategies.
@@
Temporal pattern recognition The detection of changes in arrhythmia patterns over time provides insights into disease progression and treatment efficacy. Circadian variations in arrhythmia occurrence can reveal underlying triggers, with nocturnal episodes often associated with sleep-disordered breathing and diurnal episodes linked to sympathetic activation [31, 32]. Such pattern recognition facilitates targeted therapeutic interventions and lifestyle modifications.
@@
Paroxysmal episode detection The identification of short paroxysms, particularly those lasting seconds to minutes, addresses a significant limitation of conventional monitoring systems. These brief episodes, often asymptomatic, may nonetheless contribute to stroke risk in patients with cryptogenic cerebrovascular events [34, 35]. Enhanced sensitivity for paroxysmal detection is particularly relevant for post-ablation monitoring, where early recurrence detection during the blanking period can inform subsequent management strategies.
@@
Transition state analysis The recognition of brief interruptions or transitions within arrhythmic episodes provides mechanistic insights into arrhythmia maintenance and termination [36]. Analysis of onset and termination patterns can inform catheter ablation strategies by identifying critical regions for intervention. Additionally, the detection of mode switching between different arrhythmic patterns (e.g., atrial fibrillation to atrial flutter) can reveal information about the underlying electrophysiological substrate.
Precision medicine and individualized treatment strategies These enhanced diagnostic capabilities represent a significant advancement toward precision electrophysiology, where treatment strategies are tailored to individual arrhythmia characteristics, patient physiology, and response patterns [36]. The integration of high-resolution temporal analysis with clinical risk factors enables more sophisticated risk stratification algorithms that account for both arrhythmia burden and pattern variability.
@@ 4.3 Limitations of the study
While our evaluation encompasses multiple databases and scenarios, several limitations should be acknowledged. First, our study focuses exclusively on supervised learning paradigms, which require extensive labeled data that may not always be available in clinical settings. The reliance on expert-annotated rhythm labels limits scalability to larger, unlabeled ECG datasets that are increasingly common in clinical practice. Nevertheless, the multi-epoch S4ECG design is naturally compatible with self-supervised objectives that exploit inter-epoch temporal relations for representation learning without explicit labels; future work should explore such adaptations to harness large unlabeled datasets and further improve generalization and robustness.
@@ 5 Summary and conclusion
In this work, we present S4ECG, a novel multi-epoch deep learning architecture for ECG arrhythmia detection that leverages structured state space models to capture long-range temporal dependencies. Through systematic evaluation across four major ECG databases, we demonstrate consistent and substantial improvements of multi-epoch models over conventional single-epoch approaches, with particularly notable gains in out-of-distribution scenarios. Our findings reveal that moderate temporal windows (10-20 minutes) yield optimal performance, suggesting fundamental characteristics of cardiac rhythm analysis that extend beyond dataset-specific artifacts. The enhanced robustness and generalizability of our approach address critical challenges in real-world deployment, paving the way for more accurate and reliable arrhythmia monitoring in diverse clinical settings. We envision that this work will contribute to a paradigm shift in arrhythmia detection algorithm design, encompassing both temporal modeling evolution and architectural advancement toward efficient structured state space models that better align with clinical practice and the inherent temporal nature of cardiac arrhythmias.
Acknowledgments
T.W. and N.S. conceived the study. T.W. and N.S. implemented the S4ECG architecture. T.W. conducted all experiments. T.W., W.H. and N.S. analyzed the results and T.W., W.H. and N.S. wrote the manuscript. N.S. supervised the project. All authors critically revised and approved the manuscript.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
@@ Acknowledgments
All datasets used in this study are publicly available through PhysioNet (https://physionet.org): Icentia11k, LTAFDB, AFDB, and MITDB. The source code implementation is available at the accompanying repository [37].
Appendix A Dataset details and preprocessing
A.1 Dataset details
Icentia11k The Icentia11k dataset [38, 17] serves as our primary large-scale training resource, comprising continuous single-lead ECG recordings from 11,000 patients monitored over extended periods. This dataset represents the largest publicly available collection of annotated ECG data for arrhythmia research, with recordings spanning durations from several hours to multiple weeks. Each recording is sampled at 250 Hz and includes expert annotations for various cardiac rhythm types, including atrial fibrillation (AF), atrial flutter (AFL), and normal sinus rhythm (N). The dataset’s substantial size and temporal extent make it particularly well-suited for investigating long-range temporal dependencies in cardiac rhythm analysis. Patient demographics span a diverse age range with balanced representation across gender groups, providing a robust foundation for model development.
Long-Term AF Database (LTAFDB) The LTAFDB [18] serves both as a secondary training dataset and for cross-dataset OOD evaluation, featuring 84 long-term ECG recordings from patients with documented atrial fibrillation episodes. These recordings, sampled at 128 Hz, capture the natural progression and variability of atrial fibrillation patterns over extended monitoring periods ranging from 24 to 25 hours per patient. The database includes comprehensive rhythm annotations encompassing atrial fibrillation (AF), normal sinus rhythm (N), and supraventricular tachyarrhythmia (SVTA). This dataset’s focus on atrial fibrillation provides complementary training data with different temporal characteristics compared to Icentia11k, while also enabling assessment of model performance across varied dataset scales and sampling frequencies in both ID and OOD scenarios.
MIT-BIH Atrial Fibrillation Database (AFDB) For OOD evaluation, we utilize the MIT-BIH AFDB [19], which contains 25 long-term ECG recordings specifically selected to include significant episodes of atrial fibrillation. These recordings, sampled at 250 Hz with durations of approximately 10 hours each, provide ground truth annotations for atrial fibrillation, atrial flutter, and normal rhythm segments. The database’s careful curation and well-characterized arrhythmia episodes make it an ideal benchmark for evaluating model generalization capabilities beyond the training distribution.
MIT-BIH Arrhythmia Database (MITDB) The MITDB [20] also serves as an OOD evaluation dataset, comprising 48 half-hour excerpts of two-channel ambulatory ECG recordings from 47 subjects. These recordings, sampled at 360 Hz, include comprehensive beat-by-beat annotations for various arrhythmia types. While primarily designed for beat-level classification tasks, we adapt the annotations to epoch-level rhythm classification to maintain consistency with our experimental framework. The database’s different sampling rate, recording duration, and patient population characteristics provide a stringent test of model robustness across diverse acquisition protocols and demographic variations.
A.2 Data preprocessing
We implement a standardized preprocessing pipeline that addresses the heterogeneity in sampling rates, signal durations, and annotation formats present in the original datasets. All ECG signals are standardized to a uniform sampling rate of 128 Hz using the first available channel, and rhythm annotations are processed to generate epoch-level labels through a label aggregation process that preserves the temporal distribution of rhythm types within each 30-second epoch.
Signal preprocessing The ECG signals are processed directly from the original PhysioNet format without additional normalization, as the datasets already provide calibrated recordings in millivolts (mV). This preserves the original signal characteristics and amplitude relationships as intended by the dataset creators. Given the varying native sampling rates across datasets, we standardize all ECG signals to a uniform sampling rate of 128 Hz using the resampy library, which employs high-quality resampling with automatic anti-aliasing. For LTAFDB, which already has a native sampling rate of 128 Hz, no resampling is performed. This standardization ensures computational efficiency while preserving the clinically relevant frequency components of ECG signals, which typically contain most diagnostic information below 50 Hz.
Rhythm annotation processing We extract rhythm annotations from the PhysioNet annotation files (.atr) by identifying rhythm change markers (annotated with ”+” symbols in the original datasets). These rhythm annotations define temporal segments with consistent rhythm types including normal sinus rhythm (N), atrial fibrillation (AF), atrial flutter (AFL), and supraventricular tachyarrhythmia (SVTA). Additional rhythm types such as ”Unknown” or unclassified segments are preserved in the dataset but excluded from loss calculation during training to ensure model optimization focuses on well-defined rhythm categories. We generate sample-level rhythm labels that assign each time point to its corresponding rhythm class. The rhythm label segmentation mask is then resampled to match the target sampling rate of 128 Hz, ensuring temporal alignment between signal data and rhythm annotations.
Epoch-level target generation Our implementation transforms the continuous rhythm segmentation masks into epoch-level labels through a label aggregation process. For each 30-second epoch (corresponding to 3,840 samples at 128 Hz), we count the number of samples belonging to each rhythm class within that temporal window. The epoch-level label is then computed as the fraction of time each rhythm class is present within the epoch, creating soft labels that preserve the temporal distribution of rhythm types. For example, if an epoch contains 60% normal rhythm and 40% atrial fibrillation samples, the resulting epoch label reflects these proportions rather than selecting a single predominant class. This approach maintains the rich temporal characteristics of the original PhysioNet annotations while providing epoch-level supervision suitable for multi-epoch sequence modeling. For recordings that do not divide evenly into 30-second segments, we discard the remaining partial epoch to maintain consistent input dimensions across all samples.
Appendix B Training and evaluation details
B.1 Training methodology
We split datasets at the patient level for Icentia11k and LTAFDB, and at the recording level for AFDB and MITDB. This approach ensures that no patient’s data appears in both training and evaluation sets, providing a more rigorous assessment of model robustness. For Icentia11k, we employ an 8:1:1 patient-level split, allocating 80% of patients for training, 10% for validation and model selection, and 10% for in-distribution testing. Similarly, LTAFDB follows a 3:1:1 patient-level split (60% training, 20% validation, 20% testing) to accommodate the smaller dataset size while maintaining sufficient data for each partition.
For the large-scale Icentia11k dataset, we use a streamlined training approach with 5 epochs, leveraging the substantial amount of training data (approximately 110,000 hours) to achieve convergence efficiently. For the smaller LTAFDB dataset with 84 patients and approximately 2,000 hours of recordings, we extend training to 150 epochs to ensure adequate learning from the limited data.
The optimization strategy uses the AdamW optimizer with a fixed learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-3}$ optimizing binary crossentropy as loss function with fractional targets as described above. Due to the computational demands of processing long multi-epoch sequences, we employ a memory-efficient training strategy with small batch sizes combined with gradient accumulation to achieve an effective batch size of 64. This approach enables training on sequences up to 60 epochs (30 minutes of ECG data) while maintaining computational feasibility. Models with the best macro-AUROC performance on validation are selected for final evaluation.
Multi-epoch Training Configuration. For multi-epoch models, the input size is adjusted based on the desired temporal context:
  • 1.
    2 epochs (1 min): input_size = 7,680
  • 2.
    5 epochs (2.5 min): input_size = 19,200
  • 3.
    10 epochs (5 min): input_size = 38,400
  • 4.
    20 epochs (10 min): input_size = 76,800
  • 5.
    30 epochs (15 min): input_size = 115,200
  • 6.
    40 epochs (20 min): input_size = 153,600
  • 7.
    50 epochs (25 min): input_size = 192,000
  • 8.
    60 epochs (30 min): input_size = 230,400
Each configuration maintains the fixed epoch length of 3,840 samples (30 seconds at 128 Hz), ensuring consistent epoch-level processing while varying the inter-epoch temporal context.
Appendix C Performance evaluation
We evaluate model performance using a comprehensive set of metrics tailored to the multi-label nature of our rhythm classification task. The primary evaluation metric is the macro-averaged area under the receiver operating characteristic curve (macro-AUROC), which provides equal weighting to all rhythm classes regardless of their prevalence in the dataset. This choice ensures that model performance on rare but clinically important arrhythmias (such as atrial flutter) receives appropriate consideration alongside more common rhythm types.
For each rhythm class $c$, we compute the AUROC by treating the prediction as a binary classification problem between class $c$ and all other classes. The macro-AUROC is then calculated as:
$\text{macro-AUROC}=\frac{1}{C}\sum_{c=1}^{C}\text{AUROC}_{c}$ (1)
where $C$ is the total number of rhythm classes in the dataset.
In addition to macro-AUROC, we report class-specific AUROC values to provide detailed insights into model performance for individual rhythm types. This granular analysis is particularly important for understanding model behavior across different arrhythmia types and identifying potential areas for improvement.
For OOD evaluation, models trained and selected based on Icentia11k performance are evaluated on the complete LTAFDB, AFDB, and MITDB datasets, providing comprehensive assessment of cross-dataset generalization capabilities. This evaluation strategy ensures that model selection is performed independently of the test data, preventing any form of information leakage and providing unbiased estimates of model performance across diverse clinical scenarios.
Appendix D Statistical analysis
To assess the uncertainty of model performance metrics, we provide 95% confidence intervals via empirical bootstrapping on the test set with 10,000 iterations. We report point estimates from evaluation on the complete test set and estimate confidence intervals using bootstrap resampling. Statistical significance between models is determined using bootstrap estimates of performance differences. If confidence intervals for the difference between the best-performing and other models do not include zero, the models are considered statistically significantly different at $\alpha=0.05$. During bootstrapping, samples lacking positive examples for all classes are discarded and redrawn to ensure reliable macro-AUROC computation. In result tables, we report point estimates with maximal absolute deviations between point estimates and confidence interval bounds (±values).
Appendix E Computational Details
The S4ECG model contains approximately 4.9 million trainable parameters, distributed across the hierarchical encoder-predictor architecture. All models were trained on NVIDIA A100 GPUs with 80 GB memory, leveraging the high-bandwidth memory and tensor core capabilities for efficient processing of long temporal sequences. The substantial memory requirements of multi-epoch training (particularly for 60-epoch sequences spanning 30 minutes of ECG data) necessitated the use of gradient accumulation strategies and memory-efficient implementations of the S4 architecture.
During training, we employ non-overlapping crops of the specified input size to ensure independent training samples and prevent data leakage between adjacent sequences. However, at inference time, we implement a sliding window approach to maximize the utilization of available temporal context and improve prediction robustness.
Appendix F Related work
Structured state space models Structured state space sequence (S4) models represent a recent breakthrough in sequence modeling, offering efficient alternatives to traditional recurrent and transformer architectures for long-range dependency modeling [14]. S4 models leverage the mathematical framework of state space representations to capture temporal dependencies while maintaining linear computational complexity with respect to sequence length. This efficiency makes them particularly attractive for biomedical applications involving long time series.
The theoretical foundations of S4 models enable them to capture dependencies across arbitrarily long sequences without the vanishing gradient problems that plague traditional RNNs or the quadratic complexity limitations of transformer architectures. Recent work has demonstrated the effectiveness of S4 models across diverse domains, from natural language processing to time series forecasting, establishing them as a powerful tool for sequence modeling tasks.
In the context of physiological signal analysis, S4 models offer particular advantages due to their ability to capture both short-term patterns (within individual epochs) and long-term dependencies (across multiple epochs) within a unified framework. The hierarchical application of S4 models—at both epoch-level and sequence-level processing—provides a natural fit for multi-epoch ECG analysis, enabling efficient capture of the complex temporal dependencies inherent in cardiac rhythm patterns.
Multi-epoch temporal modeling The concept of incorporating temporal context across multiple epochs has gained significant attention in biomedical signal analysis, particularly in the domain of sleep staging. The pioneering work of SeqSleepNet [39] demonstrated that processing sequences of sleep epochs jointly rather than independently leads to substantial improvements in classification accuracy and temporal consistency. The SeqSleepNet architecture employs a hierarchical approach where individual epochs are first encoded into compact representations, which are then processed by a recurrent neural network to capture inter-epoch dependencies.
This encoder-predictor paradigm has been further refined in subsequent work, with S4Sleep [13] providing a comprehensive evaluation of different architectural components for sleep stage classification. The S4Sleep study systematically investigated the impact of various deep learning architectures, including structured state space models (S4), and established design principles for multi-epoch analysis in physiological time series.
Furthermore, [24] demonstrated that moderate sequence lengths provide optimal performance, avoiding both the limited context of single-epoch models and the overfitting risks associated with excessively long sequences.
The success of multi-epoch approaches in sleep staging provides compelling motivation for their application to ECG analysis, given the similar temporal characteristics and physiological dependencies present in both domains. However, the direct translation of these approaches to arrhythmia detection requires careful consideration of the specific temporal patterns and clinical requirements inherent in cardiac rhythm analysis.
Appendix G Qualitative analysis thresholds
For the qualitative analysis presented in Figure 2 of the main text, optimal classification thresholds are determined using a false negative rate-based approach that prioritizes clinical sensitivity requirements. Specifically, for each rhythm class $i$, we compute the ROC curve and select the threshold that minimizes the absolute difference between the achieved false negative rate and a clinically acceptable target rate:
$\theta_{i}^{*}=\arg\min_{\theta}|(\text{FNR}(\theta)-\text{FNR}_{\text{target}})|$ (2)
where $\text{FNR}(\theta)=1-\text{TPR}(\theta)$ represents the false negative rate at threshold $\theta$. This approach is clinically motivated, as missing arrhythmic episodes (false negatives) are associated with higher clinical risk than false alarms in continuous monitoring scenarios [40, 41]. The method ensures that the model achieves the desired sensitivity level for detecting critical arrhythmic events, which is particularly important for life-threatening arrhythmias such as atrial fibrillation where early detection is crucial for preventing stroke and other complications.
For this study, we set $\text{FNR}_{\text{target}}=0.1$ (corresponding to 90% sensitivity) representing a high-sensitivity operating point suitable for arrhythmia screening scenarios where minimizing false negatives is prioritized.

### v2

@@ 4.2 Clinical implications
Our findings have several clinical implications. The demonstration that moderate temporal windows (10-20 minutes) yield optimal performance fundamentally challenges conventional arrhythmia detection paradigms.
@@
Furthermore, the superior out-of-distribution performance of multi-window models indicates a high degree of robustness—a critical requirement for real-world deployment, which was explicitly acknowledged for example in the FDA’s Software as a Medical Device (SaMD) guideline (Food and Administration, 2017) and the action plan on AI/ML-based SaMD (Food and Administration, 2021). This robustness addresses the well-documented challenge of domain shift in medical AI applications, where models frequently underperform when applied to patient populations, recording devices, or clinical environments that differ from training conditions. More specifically, we anticipate impact on the following use cases:
@@
Enhanced diagnostic precision and clinical burden reduction Improved overall predictive performance at fixed sensitivity can translate into higher specificity, reducing false positives, unnecessary clinical interventions, costs, and patient anxiety. False positive rates have been reported to be high in certain conventional automated systems (Bumgarner et al., 2018; Tison et al., 2018), contributing to increased emergency department utilization from consumer-grade devices (Seshadri et al., 2020; Rajakariar et al., 2020). As a result, the proposed approach can improve patient experience and resource utilization without compromising diagnostic safety.
@@
Advanced arrhythmia characterization and temporal dynamics Understanding the temporal dynamics characteristics of arrhythmia is key for more fine-grained understanding of arrhythmia subclasses, such as paroxysmal and persistent in the case of atrial fibrillation, and is believed to lead to clinically actionable insights into disease progression and treatment response (Charitos et al., 2014; De Vos et al., 2010). The improved accuracy but also temporal consistency, see Figure 3, of the proposed approach aligns with this goal.
@@
Atrial fibrillation burden assessment Precise AF burden quantification is increasingly recognized as a critical determinant of stroke risk and therapeutic decision-making (Chen et al., 2018). Studies demonstrate that even modest AF burden (>>0.5%) correlates with increased thromboembolic risk, emphasizing the clinical importance of accurate measurement. Our approach enables continuous, high-resolution burden assessment that can inform both acute and chronic management strategies.
@@
Temporal pattern recognition The detection of changes in arrhythmia patterns over time provides insights into disease progression and treatment efficacy. Circadian variations in arrhythmia occurrence can reveal underlying triggers, with nocturnal episodes often associated with sleep-disordered breathing and diurnal episodes linked to sympathetic activation (Charitos et al., 2014; De Vos et al., 2010). Such pattern recognition facilitates targeted therapeutic interventions and lifestyle modifications.
@@
Paroxysmal episode detection The identification of short paroxysms, particularly those lasting seconds to minutes, addresses a significant limitation of conventional monitoring systems. These brief episodes, often asymptomatic, may nonetheless contribute to stroke risk in patients with cryptogenic cerebrovascular events (Sanna et al., 2014; Gladstone et al., 2014). Enhanced sensitivity for paroxysmal detection is particularly relevant for post-ablation monitoring, where early recurrence detection during the blanking period can inform subsequent management strategies.
@@
Transition state analysis The recognition of brief interruptions or transitions within arrhythmic episodes provides mechanistic insights into arrhythmia maintenance and termination (Nattel and Dobrev, 2017). Analysis of onset and termination patterns can inform catheter ablation strategies by identifying critical regions for intervention. Additionally, the detection of mode switching between different arrhythmic patterns (e.g., atrial fibrillation to atrial flutter) can reveal information about the underlying electrophysiological substrate.
@@ 4.3 Limitations of the study
While our evaluation encompasses multiple databases and scenarios, several limitations should be acknowledged. First, our study focuses exclusively on supervised learning paradigms, which require extensive labeled data that may not always be available in clinical settings. The reliance on expert-annotated rhythm labels limits scalability to larger, unlabeled ECG datasets that are increasingly common in clinical practice. Nevertheless, the multi-window S4ECG design is naturally compatible with self-supervised objectives that exploit inter-window temporal relations for representation learning without explicit labels; future work should explore such adaptations to harness large unlabeled datasets and further improve generalization and robustness.
@@ 5 Summary and conclusion
In this work, we present S4ECG, a deep learning architecture that exploits structured state space models to capture diagnostically relevant long-range dependencies in electrocardiogram signals. Through systematic evaluation across four major ECG databases, we demonstrate that the joint analysis of multiple consecutive windows yields substantial improvements over conventional single-window approaches, particularly in challenging out-of-distribution scenarios. Crucially, our findings identify an optimal temporal context of 10–20 minutes, reflecting fundamental physiological timescales of cardiac rhythm dynamics that transcend dataset-specific characteristics. By significantly enhancing robustness and reducing false positive rates, this approach addresses key barriers to the reliability of automated arrhythmia monitoring. These results advocate for a methodological transition in algorithm design—moving beyond isolated morphological analysis toward efficient, temporally-aware sequence modeling that aligns with the continuous nature of cardiac pathophysiology.
Data availability statement
@@
All datasets used in this study are publicly available through PhysioNet (https://physionet.org): Icentia11k, LTAFDB, AFDB, and MITDB. The source code implementation as well as the training-validation-test data splits are available at the accompanying repository (Wang and Strodthoff, 2025b).